# RB-SCD: A New Benchmark for Semantic Change Detection of Roads and Bridges in Traffic Scenes

### v2

@@ VI Conclusion
In this work, we introduce RB-SCD dataset, a SCD dataset specifically designed for complex traffic scenarios. The dataset features fine-grained pixel-level annotations across 11 semantic change categories and one background class, encompassing a wide range of road and bridge types. It also covers diverse geographic regions across multiple countries and urban environments, making it become a comprehensive and challenging benchmarks to date for road and bridge change analysis in RS imagery. Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios. By incorporating both common and rare semantic transformations, RB-SCD provides a valuable resource for the development and evaluation of CD models in traffic contexts.
To address the unique challenges presented by this dataset, we propose a novel multimodal CD framework, MFDCD, which leverages both spatial and frequency domain representations to enhance feature learning. By introducing frequency-domain analysis into the CD pipeline, MFDCD captures subtle semantic differences and structural changes that may be overlooked by conventional spatial-domain approaches. Our method also incorporates an effective fusion mechanism to integrate multimodal features, enabling more accurate detection of complex changes such as occluded road regions. Extensive experiments conducted on the RB-SCD dataset and public CD datasets demonstrate the superior performance of MFDCD across various metrics when compared to state-of-the-art CD methods. The results highlight the robustness, generalization ability, and semantic discrimination power of our approach in both straightforward and highly ambiguous change scenarios.
In future work, we plan to extend our exploration of road and bridge change detection by integrating additional modalities such as elevation data or temporal sequences, and further investigating the role of frequency-domain representations in improving the interpretability and adaptability of CD models. As the label imbalance analysis and challenges for RB-SCD dataset discussed in Sec. III-B, we also plan to expand the scale of the dataset and enrich it with more detailed spatiotemporal annotations to support broader research tasks such as urban infrastructure evolution and intelligent transportation systems. Furthermore, while MFDCD achieves strong performance, it is not yet lightweight enough for deployment in resource-constrained environments. Future research will focus on model compression and efficient architecture design to improve its applicability in real-time and edge scenarios.

### v3

@@ 6 Conclusion
In this work, we introduce RB-SCD, a SCD dataset specifically designed for complex traffic scenarios. RB-SCD provides comprehensive fine-grained pixel-level annotations across 11 semantic change categories, encompassing a wide range of road and bridge types. It also covers diverse geographic regions across multiple countries and urban environments, making it a challenging benchmark for road and bridge change analysis in traffic scenes. Notably, RB-SCD supports both BCD and SCD tasks, offering flexibility for different application scenarios. We believe this benchmark represents an important first step towards advancing SCD of roads and bridges, and we envision future extensions to larger-scale and more balanced datasets.
In addition, we propose the MFDCD framework, which pioneers a new paradigm by integrating multimodal features in the frequency domain. It tackles the core challenges with two specialized components: the DFC module, which leverages wavelet analysis to robustly model the evolution of linear structures, and the TFF module, which acts as a semantic reasoning engine to disambiguate visually similar land covers. Extensive experiments on RB-SCD and three public datasets validate the superiority of MFDCD over state-of-the-art methods, highlighting its robustness, generalization, and strong discrimination power.
Limitation and Feature Work. Nevertheless, we acknowledge its limitations. The performance of MFDCD still degrades under extreme shadow or for rare, under-represented change classes, and the framework is not yet lightweight enough for deployment in resource-constrained scenarios. Future work will therefore focus on three directions: (1) expanding RB-SCD with more rare-class samples and temporal sequences; (2) extending our frequency-driven paradigm to topology-aware change reasoning for better connectivity analysis and pseudo-intersection recognition; and (3) developing lightweight model variants via compression and efficient architecture design to enhance applicability in real-time and edge-computing scenarios.
Acknowledgement
This work was supported in part by the NSFC Key Project of Joint Fund for Enterprise Innovation and Development under Grant U24A20342, and in part by the National Natural Science Foundation of China under Grant 62576006 and 61976004.